# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ 4.3. Strong convergence
    There exists $C\geq 0$ such that for all $(t_{1},x_{1},a_{1}),(t_{2},x_{2},a_{2})\in[0,T]\times\mathbb{R}^{d}\times A$, $u\in E$,
    $\left|b(t_{1},x_{1},a_{1})-b(t_{2},x_{2},a_{2})\right|+d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|+d_{A}(a_{1},a_{2})\right),$ (4.17)
    and for all $x,y\in\mathbb{R}^{d}$ and $t\in[0,T]$,
    $\left|\sigma(t,x)-\sigma(t,y)\right|\leq C\left|x-y\right|.$ (4.18)
@@
    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential  [^3] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$..
@@
If $\bm{\pi}(\cdot|t,x)$ satisfies (4.7) with $\Sigma(t,x)=\Sigma$, then we can write $a=\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$. In this case, the assumptions for $\phi$ in (H.6) hold. This corresponds to the commonly used Gaussian policies considered in [32] and many others.
@@
Suppose (H.2) and (H.6) hold. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
@@
The proof of Theorem 4.4 is given in Section 7.1.4. It involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval and subsequently controlling the overall error using concentration inequalities for sub-exponential martingale difference sequences.
@@
Unfortunately, such a strong convergence fails when the volatility is controlled, as shown in the following example.
@@
$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-Q^{1/2}\right)\mathrm{d}W_{s}.$ (4.21)
Note that $W$ is a Brownian motion under the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ with $\mathcal{F}_{t}:=\sigma\{a_{\delta(s)}:s\leq t\}\lor\mathcal{F}^{W}_{t}$. It follows from Itô’s isometry that
$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-Q^{1/2})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-Q^{1/2})^{2}\right],$ (4.22)
@@ 5.1. Accuracy of the Monte-Carlo evaluation
The agent can only discretely sample the actions at the pre-specified time grid and observe the state variable at the same time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$. Therefore, the agent can sample the process $X^{\mathscr{G}}_{s}$. For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ [^4] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward need to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme. Under suitable assumptions on the regularity conditions on the function $r$, the order of our analysis does not change. (cf. [18]) Hence, the expected reward collected by the agent is
@@
In, the expected reward with a continuous-time execution of the policy is given by (2.6):
@@
[…]
@@
Assume (H.1) and (H.2) hold. If (H.5) and the conditions in Proposition 5.1(1) hold, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
@@ 5.2. Accuracy of Temporal-Difference learning
In particular, suppose $V^{\theta}$ is a family of parameterized functions of $(t,x)$ with $\theta$ a finite-dimensional parameter. TD learning concerns the increment in the form “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward “$r(t,\widetilde{X}_{t})\mathrm{d}t$”. In the continuous-time framework, these two increments have different natures. “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter “$r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time. In theory, [19] identify the continuous-time counterpart of TD learning as solving the martingale orthogonality condition. More precisely, if $\theta^{*}$ is the root to
@@ 5.3. Accuracy of policy gradient
The policy gradient is one of the most commonly used methods in policy optimization in RL. The well-celebrated (model-free) policy gradient representation (e.g., [31]) gives an unbiased estimator for the gradient of the value function with respect to certain parameters in the policy. For example, the stochastic policy $\bm{\pi}^{\psi}$ is parameterized by $\psi$, a finite-dimensional parameter, and it admits a density function with respect to the Lebesgue measure. We denote the density function by $H^{\psi}(t,x,a)=\frac{\bm{\pi}^{\psi}(\mathrm{d}a|t,x)}{\mu(\mathrm{d}a)}$. Then we know the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$. The policy gradient concerns about the derivative of $\tilde{J}(t,x;\bm{\pi}^{\psi})$ with respect to $\psi$, i.e., $\frac{\partial\tilde{J}(t,x;\bm{\pi}^{\psi})}{\partial\psi}$. For simplicity, we focus on the initial time-state gradient and suppress the time-state pair’s dependence.
@@
where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $[t_{i},t_{i+1}]$.
@@ 5.4. Accuracy of risk-sensitive q-learning
6. Conclusion
This paper addresses a fundamental problem in the continuous-time RL framework proposed in [36] by answering the relation between the aggregated dynamics (referred to as the “exploratory dynamics” therein) and the actual sampling process. In terms of SDE theory, we provide a rigorous treatment of the “law-of-large-numbers” heuristics used in [36] to demonstrate how the original dynamics lead to the aggregated dynamics. Our analysis is particularly crucial for understanding how discrete sampling relates to the continuous-time RL problems. We establish the convergence and convergence rate of discretely sampled stochastic policies, i.e., the sampled dynamics, to the aggregated dynamics in different senses. Furthermore, we illustrate the bias and variance of the estimate for several important quantities in analyzing continuous-time RL algorithms when using sampled dynamics, including values, temporal differences, policy gradients, and quadratic variations. These bias and variance analyses are critical for the convergence of stochastic approximation algorithms, and they provide guidance on properly tuning the interaction frequency to foster faster convergence of RL algorithms.
The results of this paper also stimulate more interesting open questions for future research. For example, other sampling methods that could achieve higher accuracy in the RL context, any exact simulation methods for controlled SDEs under stochastic policies if the coefficients are all known/unknown, and a full account for the convergence of the gradient-based continuous-time RL algorithm.
@@ 7. Proofs of Main Results
Proof.
@@
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.1)
Let $v_{f}(t,x;t_{j})$ be the solution of (7.1). For simplicity, we dorp the $f$ and $t_{j}$ and use the notation $v(t,x)$. By (H. 3), $v(\cdot,\cdot)$ is smooth and $\|v(\cdot,\cdot)\|_{C_{p}^{4}}<+\infty$.
Noting that if $v$ is the solution of (7.1), then by the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
@@
It suffices to estimate every $\mathbb{E}[e_{i}]$. For simplicity, denote $\partial_{t}:=\partial/\partial t$ and $\partial_{x}:=\partial/\partial x$. It follows from Itô’s lemma (see the dynamic (3.3) of $X^{\mathscr{G}}$) that
@@
with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where in the second equality, we use the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1). Noting that $\xi_{i-1}$ is independent of $X^{\mathscr{G}}_{t_{i-1}}$, we have
@@
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}=0,$
where the first equality follows form taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (7.1). Therefore,

### v2

@@ 4.3. Strong convergence
    There exists $C\geq 0$ such that for all $x,y\in\mathbb{R}^{d}$, $a_{1},a_{2}\in A$, and $t\in[0,T]$,
    $\left|b(t,x,a_{1})-b(t,y,a_{2})\right|+\left|\sigma(t,x)-\sigma(t,y)\right|$ $\leq C\left(\left|x-y\right|+d_{A}(a_{1},a_{2})\right),$
    $\left|b(t,0,0)\right|+\left|\sigma(t,0)\right|$ $\leq C.$
@@
    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential. [^5] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$. Moreover, there exists $C\geq 0$ such that for all $(t_{1},x_{1}),(t_{2},x_{2})\in[0,T]\times\mathbb{R}^{d}$, and $u\in E$,
    $d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|\right).$ (4.19)
Assumption (H.62) holds for commonly used Gaussian policies considered in [37] and many others.
@@
Let $A={\mathbb{R}}^{k}$, and $\bm{\pi}(\cdot|t,x)$ be a Gaussian policy of the form (see e.g., [37]):
${\bm{\pi}(\mathrm{d}a|t,x)}=\varphi(a|\mu(t,x),\Sigma^{2})\mathrm{d}a,$ (4.20)
where $\varphi(\cdot|\mu,\Sigma^{2})$ denotes the Gaussian density with mean $\mu$ and variance $\Sigma^{2}$. In this case, one can define the sampling function $\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$, with a standard normal random variable $\xi$. If $\mu$ is Lipschitz continuous in $x$ and $1/2$-Hölder continuous in $t$, then $\phi$ satisfies (H.62).
More generally, Assumption (H.62) holds for stochastic policies with sufficiently regular densities.
Example 4.6.
Let $A=[\alpha,\beta]$ for $\alpha<\beta$, and $\bm{\pi}({\mathrm{d}}a|t,x)=h(t,x,a){\mathrm{d}}a$ with a continuous density $h:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$. In this case, by the inverse transform method, one can sample from $\bm{\pi}({\mathrm{d}}a|t,x)$ by $\phi(t,x,\xi)$, where $\xi$ is a uniform random variable on $[0,1]$, and $0,1]\ni u\mapsto\phi(t,x,u)\in A$ is the inverse function of the map $A\ni r\mapsto H(t,x,r)\coloneqq\int_{\alpha}^{r}h(t,x,a){\mathrm{d}}a\in[0,1]$. If there exists $C,c>0$ such that for all $(t,x),(t^{\prime},x^{\prime})\in[0,T]\times{\mathbb{R}}^{d}$ and $a\in A$,
$h(t,x,a)\geq c,\quad|h(t,x,a)-h(t^{\prime},x^{\prime},a)|\leq C(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|),$ (4.21)
then $\phi$ satisfies (H.62).
Indeed, $|\phi(t_{0},x_{0},\xi)|$ is sub-exponential since $A$ is bounded. Moreover, since $\phi$ is the inverse of $H$,
$H(t,x,\phi(t,x,u))-H(t,x,\phi(t^{\prime},x^{\prime},u))+H(t,x,\phi(t^{\prime},x^{\prime},u))-H(t^{\prime},x^{\prime},\phi(t^{\prime},x^{\prime},u))$
$=H(t,x,\phi(t,x,u))-H(t^{\prime},x^{\prime},\phi(t^{\prime},x^{\prime},u))=u-u=0,$
which along with the mean value theorem and the definition of $H$ yields that
$\phi(t,x,u)-\phi(t^{\prime},x^{\prime},u)=\frac{\int_{\alpha}^{\phi(t^{\prime},x^{\prime},u)}(h(t^{\prime},x^{\prime},a)-h(t,x,a)){\mathrm{d}}a}{\partial_{r}H(t,x,\tilde{a})}\leq\frac{\int_{\alpha}^{\beta}|h(t^{\prime},x^{\prime},a)-h(t,x,a)|{\mathrm{d}}a}{h(t,x,\tilde{a})}$
for some $\tilde{a}$ between $\phi(t,x,u)$ and $\phi(t^{\prime},x^{\prime},u)$. Using (4.21) shows that $\phi$ satisfies (4.19).
Similar arguments extend to multidimensional action sets $A$ using the conditional inverse sampling method (see e.g., [8, Chapter 6.3]).
Now we prove the strong convergence of the sampled dynamics under Assumption (H.6). Note that (H.6) implies (H.1) and (H.2), which ensure that the aggregated state process $\widetilde{X}$ and the sampled state process $X^{\mathscr{G}}$ are $p$-th power integrable for all $p\geq 2$ (see Lemma 6.5).
@@
Suppose (H.6) holds. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
@@
Proof.
See Section 6.1.4. ∎
The proof of Theorem 4.4 is given in Section 6.1.4. The analysis involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval by conditioning on the realization of the Brownian motion, and then controlling the overall error using concentration inequalities for conditional sub-exponential martingale difference sequences.
@@
However, such a pathwise convergence fails when the volatility is controlled, as shown in the following example.
@@
$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-\sqrt{Q}\right)\mathrm{d}W_{s}.$ (4.24)
Note that $W$ is a Brownian motion under the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ with $\mathcal{F}_{t}:=\sigma\{a_{\delta(s)}:s\leq t\}\lor\mathcal{F}^{W}_{t}$. By Itô’s isometry and Fubini’s theorem,
$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-\sqrt{Q})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-\sqrt{Q})^{2}\right],$ (4.25)
@@
4.4. Numerical illustration
We present numerical results illustrating the theoretical strong and weak convergence rates for two representative one-dimensional cases: uncontrolled volatility $\mathrm{d}X_{t}=a_{t}\mathrm{d}t+\mathrm{d}W_{t}$ and controlled volatility $\mathrm{d}X_{t}=a_{t}\mathrm{d}W_{t}$. In both cases, we set the time horizon $[0,1]$ and the initial state $X_{0}=0$, and consider sampling actions from a standard Gaussian policy (i.e., $a_{t}\sim\mathcal{N}(0,1)$) over a uniform grid $\mathscr{G}=\{i\frac{T}{n}\}_{i=0}^{n}$, where $n\in{\mathbb{N}}$ is the number of grid points. We consider the quantity $|\mathbb{E}[(X^{\mathscr{G}}_{T})^{4}]-\mathbb{E}[(\widetilde{X}_{T})^{4}]|$ to examine the weak convergence of the sampled dynamics, and the root mean squared error (RMSE) $\sqrt{\mathbb{E}[|X^{\mathscr{G}}_{T}-\widetilde{X}_{T}|^{2}]}$ to assess the strong convergence. All expectations are estimated using the Monte Carlo method with 50000 samples. Each point in the following plots is an average over 100 independent runs, and the shaded regions represent the 95% confidence intervals, estimated using 1000 bootstrap resamples.
Figure 1. Weak and strong convergence analysis for the uncontrolled volatility case $\mathrm{d}X_{t}=a_{t}\mathrm{d}t+\mathrm{d}W_{t}$ with actions $a_{t}\sim\mathcal{N}(0,1)$. The corresponding aggregated dynamics is $\mathrm{d}\tilde{X}_{t}=\mathrm{d}W_{t}$. The test function is $f(x)=x^{4}$ and $T=1$. Left: Weak error versus the number of grid points $n$. Right: RMSE versus the number of grid points $n$. Both axes are on log scales.
Figure 1 presents the convergence results for the case with uncontrolled volatility. The left panel shows that the weak error decays linearly with the sampling frequency on a log-log scale, with a slope of approximately 1-1. This confirms the first-order convergence rate established in Theorem 4.1. The right panel shows that the RMSE also decays linearly on the log-log plot, with a slope of approximately 1/2-1/2, confirming the strong convergence rate of $O(|\mathscr{G}|^{1/2})$ as stated in Theorem 4.4. The confidence interval for the strong error is not visible due to the very small variance of the estimator.
Figure 2. Weak and strong convergence analysis for the controlled volatility case $\mathrm{d}X_{t}=a_{t}\mathrm{d}W_{t}$ with actions $a_{t}\sim\mathcal{N}(0,1)$. The corresponding aggregated dynamics is $\mathrm{d}\tilde{X}_{t}=\mathrm{d}W_{t}$. The test function is $f(x)=x^{4}$ and $T=1$. Left: Weak error versus the number of grid points $n$. Right: RMSE versus the number of grid points $n$. Both axes are on log scales.
Figure 2 presents the convergence results for the second case, where the volatility is controlled. Similar to the first case with uncontrolled volatility, the left panel shows that the weak error exhibits a first-order decay with a slope of 1-1, consistent with Theorem 4.1. In contrast, the right panel, which depicts strong convergence, shows that the RMSE remains essentially constant as the number of grid points $n$ increases. This provides a clear numerical illustration of the failure of strong convergence under controlled volatility, confirming the theoretical argument presented in Example 4.7.
@@ 5.1. Accuracy of the Monte-Carlo evaluation
The agent can only discretely sample the actions at the pre-specified time grid and observe the state variable at the same time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$. Therefore, the agent can sample the process $X^{\mathscr{G}}_{s}$. For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$. [^6] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward needs to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme. Under suitable assumptions on the regularity conditions on the function $r$, the order of our analysis does not change. (cf. [20]) Hence, the expected reward collected by the agent is
@@
In contrast, the expected reward with a continuous-time execution of the policy is given by (2.6):
@@
    If (H.4) holds, and $r\in\mathcal{H}_{T}^{(l)},\tilde{r}\in\mathcal{H}_{T}^{(l)},h\in\mathcal{H}^{(l+2)}$ and $\tilde{r}(t,\cdot)\in\mathcal{H}^{(l+2)},\ 0\leq t\leq T$, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
@@
Assume (H.1) and (H.2) hold. If (H.5) and the conditions in Proposition 5.1(1) hold, and $\sup_{a\in A}\|r(\cdot,\cdot,a)\|_{C_{p}^{2}}<+\infty$, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\sup_{a\in A}\|r(\cdot,\cdot,a)\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
@@ 5.2. Accuracy of Temporal-Difference learning
In particular, suppose $V^{\theta}$ is a family of parameterized functions of $(t,x)$ with $\theta$ a finite-dimensional parameter. TD learning concerns the increment in the form $\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward $r(t,\widetilde{X}_{t})\mathrm{d}t$”. In the continuous-time framework, these two increments have different natures. $\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter $r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time. In theory, [21] identify the continuous-time counterpart of TD learning as solving the martingale orthogonality condition. More precisely, if $\theta^{*}$ is the root to
@@ 5.3. Accuracy of policy gradient
The policy gradient is one of the most commonly used methods in policy optimization in RL. The well-celebrated (model-free) policy gradient representation (e.g., [36]) gives an unbiased estimator for the gradient of the value function with respect to certain parameters in the policy. For example, the stochastic policy $\bm{\pi}^{\psi}$ is parameterized by $\psi$, a finite-dimensional parameter, and it admits a density function with respect to the Lebesgue measure. We denote the density function by $H^{\psi}(t,x,a)=\frac{\bm{\pi}^{\psi}(\mathrm{d}a|t,x)}{\mu(\mathrm{d}a)}$. Then we know that the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$. The policy gradient concerns about the derivative of $\tilde{J}(t,x;\bm{\pi}^{\psi})$ with respect to $\psi$, i.e., $\frac{\partial\tilde{J}(t,x;\bm{\pi}^{\psi})}{\partial\psi}$. For simplicity, we focus on the initial time-state gradient and suppress the time-state pair’s dependence.
@@
where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $t_{i},t_{i+1}]$.
@@ 6. Proofs of Main Results
The following lemma can be obtained by using Itô’s formula and Lemma 3.1. The proof is given in Appendix A.
Lemma 6.1.
Suppose (H.1) holds, and $g\in C_{p}^{2,0}$. Then there is a constant $C\geq 0$ such that for all grids $\mathscr{G}$ and $s\in[0,T]$,
$\left|\mathbb{E}\left[g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})-g(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})\right]\right|\leq C\|g\|_{C_{p}^{2,0}}|\mathscr{G}|.$ (6.1)
Based on Lemma 6.1, we prove Theorem 4.1.
Proof of Theorem 4.1.
@@
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d},$ (6.2)
where $\mathcal{L}$ is the generator of the aggregated dynamic $\widetilde{X}$ (2.4) and is defined by (4.3). Denote by $v:[0,t_{f}]\to{\mathbb{R}}$ the solution of (6.2). By (H.3), $v$ is smooth and $\|v\|_{C_{p}^{4}}<+\infty$.
By the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$, and hence
@@
It suffices to estimate each $\mathbb{E}[e_{i}]$. For simplicity, denote $\partial_{t}:=\partial/\partial t$ and $\partial_{x}:=\partial/\partial x$. It follows from Itô’s lemma (see the dynamic (3.3) of $X^{\mathscr{G}}$) that
@@
with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where the second equality used the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1). Noting that $\xi_{i-1}$ is independent of $X^{\mathscr{G}}_{t_{i-1}}$, we have
@@
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]{\color[rgb]{0,0,0}=0},$
where the first equality follows from taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (6.2). Therefore,